How many quarts of pure antifreeze should Frida add to 10 quarts of a 20% antifreeze solution to make a 50% antifreeze solution?
Answer by checkley71(8403)

X+10*.2=(X+10).5
X+2=.5X+5
X-.5X=5-2
.5X=3
X=3/.5
X=6 ANSWER FOR THE AMOUNT OF PURE ANTIFREEZE NEEDED.
PROOF
6+2=16*.5
8=8